### v1

Abstract
The reawakened era of lunar exploration is defined by a strategic shift from temporary visits to a sustained international and commercial presence, resulting in an unprecedented demand for a robust and continuously available communication infrastructure. The conventional direct-to-Earth communication architecture relies on limited and oversubscribed deep space networks, which are further challenged by the radiative environment and insufficient visibility in certain areas of the cislunar domain. We address these issues by proposing a foundational move toward inter-domain space network cooperation by introducing architectures based on near space networks. They can directly service lunar surface users or, via cislunar relays, by forming a resilient and multi-layered communication backbone. First, we establish a unified link analysis framework incorporating frequently disregarded environmental factors, such as the Moon’s variable illumination, to provide a high-fidelity performance evaluation. Second, we assess architectures’ reliability based on the outage risk, essential for quantifying the operational robustness of communication links. Finally, to manage the inherent dynamism of architectures, we propose an inter-domain space digital twin—an dynamic decision-making engine that performs real-time analysis to autonomously select the best communication path, ensuring high and stable reliability while simultaneously optimizing power consumption. Overall, our paper provides a holistic architectural and conceptual management framework, emphasizing the necessity of lunar communications to support a permanent human and economic foothold on the Moon.
Lunar communication, cislunar space, deep space, Moon, near space, networks of networks, relaying.
1 INTRODUCTION
The 21st century marks a new era of lunar exploration by shifting from temporary visits to establishing a permanent and sustainable presence on the Moon [1]. A diverse ecosystem of commercial companies and international partners affiliating with long-term ambitions for settlement and resource utilization are all part of this revitalized global endeavor [2]. This strategic shift from transient missions to sustained presence necessitates continuous operations, a multitude of interacting assets [3], including habitats and rovers [4], and sophisticated logistical chains [5]. Therefore, a continuous and reliable communication backbone connecting the Moon to Earth without any interruptions is a vital requirement for these missions, as it has a direct impact on each phase of the mission [6].
The conventional way of lunar communication is to establish a direct-to-Earth (DTE) link, which is advantageous due to its simplicity and lower latency [7]. Despite its simplicity, this solution is accompanied by limitations and challenges that cannot be overlooked, making the conventional communication architecture mismatched to the demands of the new lunar era. The recent attempts by Japan’s iSpace highlight these factors that led to the loss of contact with the landers and their eventual crash on the Moon’s surface [8], despite both missions being close to achieving a soft landing. Particularly, the visibility of certain areas on the Moon from a ground station significantly varies during a day, resulting in limited line-of-sight (LoS) availability [9]. To mitigate this, NASA operates three deep space network (DSN) facilities strategically located about apart, but even with this global distribution, any single location on the near side of the Moon has limited and intermittent contact windows [10]. This problem is more critical for the lunar far side, which is permanently shielded from Earth, making DTE communication unfeasible without a relay [11].
Establishing a DSN facility is also not economically scalable. To overcome the immense path loss due to the distance, ground stations must be equipped with advanced systems, such as giant antennas [12] (such as those with diameters of 70 meters) and high-power amplifiers that transmit signals in the kilowatt range [13]. This implies that constructing and operating a DSN station can incur costs in the hundreds of millions of dollars, which is why increasing the number of DSN stations for direct communication is not an easily applicable option. Additionally, the DSN is a heavily oversubscribed resource and stations are not always used for lunar missions, as they are essentially intended to support deep space missions, creating regulatory constraints [14] and scheduling conflicts [15]. Therefore, access to DTE links is challenging for non-agency users such as commercial companies or academic institutions. As these challenges become more apparent with the increasing number of missions, cislunar space networks (CSNs) have been planned by both commercial and national actors to reduce costs, expand access, and ensure reliable communication [16].
This new paradigm shifts away from centrally managed point to point links to a flexible [17], scalable [18], and interoperable [19] network of networks such as NASA’s emerging framework, LunaNet [20]. These solutions are designed to provide the robust and on-demand connectivity required to sustain a permanent presence on the lunar surface by burgeoning multi-layered constellations of relay satellites [21] and advanced networking protocols [22]. CSN architectures in orbits such as low lunar orbit (LLO) [23] or near rectilinear halo orbit are proposed to overcome connectivity problems at the Moon [24]. However, while CSNs ensure local coverage even for assets on the far side or in shadowed polar regions, a critical architectural question remains: how do these networks create a robust and efficient transmission back to Earth? This stimulates inter-domain space networks by facilitating state-of-the-art cooperation among multi-layered CSNs and near space networks (NSNs).
To advance this cooperation between CSNs and NSNs, a more versatile but rational analysis framework is required, given that both cislunar and near space domains111Near space domain is the operational region extending from the Karman line [25] to geostationary orbit (GEO) at km while the cislunar space domain is the volume from GEO to approximately km beyond the Moon’s second Lagrange point [3]. have different operational considerations, such as intense traffic in near space [26] or quite limited sources in cislunar space [27]. The high congestion of NSNs, especially at lower orbits, and related problems [28] make the concept of a GEO network with two-way communication links more reasonable for inter-domain space networks, as they also provide more stable and wide coverage in both directions. However, the need for multifaceted and generalized link analysis models still exists, while current models are often insufficient because their oversimplified structure fails to capture the unique aspects of each domain and the complex interactions between them [2]. A critical oversight in many models is the impact of variable thermal noise [29], where the changing brightness temperature of the lunar surface directly contributes to the system noise temperature of a receiver [30] and impacts link reliability by degrading the signal-to-noise ratio (SNR), especially at higher frequency bands like Ka-band [31, 32, 33]. This emphasizes the need of a unified link analysis framework for inter-domain space communication, which embodies these refined environmental effects to provide an accurate and holistic architectural assessment.
The holistic assessments to quantify and mitigate the risk of communication failure under dynamic real-time conditions are mandatory [34], as a link failure for high stakes operations, especially those involving crewed missions or expensive robotic assets, is not an acceptable option. This situation elevates the importance of rigorous reliability assessment and motivates the outage probability as a critical performance metric to analyze the operational robustness and trustworthiness of critical systems [35]. Furthermore, the performance of any given communication architecture—whether DTE, LLO relay, or a hybrid LLO-GEO system—is not always optimal due to the dynamic interplay of orbital mechanics, network topology, and environmental conditions throughout the missions. Therefore, relying solely on architectural designs for inter-domain space networks is not entirely satisfactory, which raises a compelling need for dynamic management of these architectures. Such a system must dynamically select the best communication path—switching between lunar communication architectures—to prioritize link reliability while also accounting for domain-level operational limitations like power consumption [36]. We therefore propose an inter-domain space digital twin. Its decision-making engine uses a high-fidelity virtual model of the network, updated with real-time telemetry, to predict performance, diagnose faults, and autonomously command the network to adapt its routing configuration, ensuring the most reliable and efficient communication path is always active.
In this paper, we propose to advance lunar communication first through inter-domain space networks by introducing and comparing architectures with practical and dynamical models. We then aim to enhance the utilization efficacy of the architectures with intelligent management. Fig. 1 shows our integrated approach from physical assets to their dynamic orchestration. Our main contributions are listed as follows.
We addresses the severe limitations of conventional lunar communication, which hinder continuous access and mission scalability, by proposing a comprehensive collaboration of inter-domain space networks. We introduce two architectures that integrate a GEO network (in the near space domain) with a LLO network (in the cislunar space domain), enabling more robust and cost-effective communication essential for future lunar endeavors.
We present a unified link analysis framework that holistically integrates critical link budget factors by incorporating unique aspects of near and cislunar space. Particularly, our framework includes the transmission dynamics and propagation mechanisms among nodes and a refined characterization of thermal noise sources, such as the variable brightness temperature of the lunar surface (a distinction often overlooked in generalized models). This comprehensive approach provides the essential foundation for precise performance assessment of diverse inter-domain lunar communication architectures.
To ensure mission success, we extend our link analysis with a reliability assessment based on outage risk, addressing the paramount need for continuous communication. We formulate the outage probability for introduced architectures to provide vital insights into their operational robustness and trustworthiness under dynamic conditions where communication failures can have severe consequences.
Overcoming inherent trade-offs in lunar communication architectures and ensuring seamless, high performance communication, we introduce an inter-domain space digital twin. This decision-making engine moves beyond static network designs by offering a dynamic optimization approach that continuously assesses real-time link conditions. It dynamically selects the best scenario by prioritizing reliability while also optimizing power utilization, which is critical for managing the complexity of emerging lunar architectures and enabling autonomous and robust missions.
We validate our contributions with presented numerical results that are obtained with the simulations in a well-known digital mission engineering tool, Satellite Tool Kit (STK) [37]. The analysis of in-orbit dynamics across four communication scenarios reveals significant fluctuations in link geometry and visibility, highlighting the inherent unreliability of each architecture. Reliability assessments, based on outage probability, quantify the performance of each architecture, demonstrating that while introduced NSN integrated architectures usually offer advantages, no single option provides continuous, robust connectivity under all dynamic conditions. The proposed inter-domain space digital twin successfully maintains a stable and highly reliable communication link, even under challenging and variable noise environments, by dynamically selecting the best scenario based on real-time link conditions and expected reliability targets.
2 UNIFIED FRAMEWORK FOR LINK ANALYSIS
To assess the suggested lunar communication architectures, we need a detailed analysis system that can represent the complicated and changing environment at different space domains. Conventional link models are frequently insufficient, as their simple structure fails to represent the distinctive features of the cislunar and near space realms. This section, therefore, develops a unified link analysis framework that holistically integrates critical link budget factors to offer accurate analysis. This comprehensive approach provides the essential foundation for the precise performance and reliability assessments conducted in the subsequent sections.
2.1 Received Signal Power
The transmitter antenna sends the communication signal to the receiver antenna at the observation planes . The received signal power as a function of distance () between the transmitter and the receiver is given below:
where is the transmit power, is the path loss and and denote the antenna power gain at the transmitter and the receiver, respectively. The atmospheric loss due to the gaseous attenuation, is included for links with terrestrial networks on Earth. and are the losses due to transmitter and receiver equipment, respectively.
2.1.1 Path Loss
The locations of a transmitter and receiver are defined as and in terms of Earth-Centered cartesian coordinates. The distance between the transmitter and receiver is calculated by considering LoS as follows
The free space path loss is given as
where denotes the wavelength.
2.1.2 Antenna Power Gain
The power gain of an antenna, is calculated for the corresponding elevation and azimuth angles at the observation planes as
Here, is the radiation efficiency111 also known as the ohmic efficiency factor or the ohmic-loss factor since it originates from the ohmic-losses. In the case of no antenna losses . and is the directivity gain as
where denotes the antenna power pattern.
2.1.3 Atmospheric Loss
The atmospheric loss () is calculated by considering the gaseous attenuation in accordance with the recommendations by the international telecommunication union as below [38],
is total water vapour attenuation and is the Earth centered elevation angle. is total oxygen attenuation where represents the equivalent height attributable to the oxygen and is specific attenuation because of dry air oxygen, nitrogen and Debye attenuation. Note that is not included by setting dB in analyses for the links where the receiver is not on the ground.
2.2 Thermal Noise
Thermal noise arises from the thermal emissions of electronic devices. It is superimposed on the transmitted signal at the receiver and typically modeled with additive white Gaussian noise (AWGN). To determine SNR, we must characterize the thermal noise sources, which are collectively represented by an operational equivalent noise temperature . The total noise power is given as follows [29, 30]
where is the Boltzmann constant and is the bandwidth. combining all noise contributions is computed as
where is the sky noise temperature, is the receiver noise temperature, is the transmission line temperature, is internal antenna noise temperature and is the additional external antenna noise temperature. shows the thermal efficiency of the transmission line, scaling the noise contributions appropriately for the antenna aperture reference point. Their calculations are detailed as follows.
2.2.1 Sky Noise Temperature
originates from the cosmic microwave background and atmosphere as
where is the physical temperature of the atmosphere and is cosmic microwave background noise.
2.2.2 Antenna Noise Temperature
The external antenna noise temperature, is aroused from the brightness temperature of the subtended body and calculated [39] as follows
Here, is the antenna beam solid angle [40] as
where . denotes the normalized antenna power pattern. In this paper, it is defined for the antennas with the diameter [41] as below
The equation (10) is simplified as by assuming that there is no other radiation sources in antenna surrounding and thus the brightness temperature does not change. Then, the increase in as a function of the brightness temperature is calculated for antenna pointed to the bright body as
If the antenna is pointed at the lunar surface, the correction factor is added in above equation (13), and it is written as follows:
The reason for this correction is that the noise sources in radio astronomy exhibit random polarization, which causes the antenna to receive only half of the power [40]. The solid angle at the distance subtended by the bright body (Moon or Earth) is for any in steradian as follows
where denotes the radius of the bright body that is km and km for Moon and Earth, respectively. , the internal antenna noise temperature depends on the physical temperature of the antenna and is calculated as below
2.2.3 Transmission Line Temperature
The transmission line noise temperature depending on the physical temperature of the transmission line () is as follows
2.3 Signal and Channel Models
The calculation of the received signal and noise power requires the consideration of different factors through different architectural scenarios depending on the source and destination: Moon (M), LLO, GEO, Earth (E). For the scenario , the received signal between the source and the destination is defined as follows
Here, represents the transmitted symbol with and is the channel fading coefficient between the source and destination nodes. is a complex AWGN following . The instantaneous SNR for each link is computed as
All links between a transmitter on the lunar surface to a receiver at any location are subject to Rician fading [42] and the cumulative distribution function of the SNR is as follows:
where and represent the Rician fading parameter and the average received SNR.
3 RELIABILITY ASSESSMENT FOR LUNAR COMMUNICATION ARCHITECTURES
Building upon the link analysis framework, this section quantifies the reliability of lunar communication architectures for inter-domain space network cooperation. We utilize the outage probability for each network architecture as the reliability performance metric since failures or substantial deterioration in communication quality might compromise mission objectives and even threaten assets or human lives. For any given communication link of -th scenario is given depending on the minimum SNR threshold () and instantaneous SNR () as [43]
The outage probability for the first scenario where the information is transmitted directly from Moon-to-Earth is formulated as
For the second scenario, communication occurs from the Moon to a LLO relay, and then from the LLO relay to Earth ground station. As a failure of either the LLO-to-Earth link or the Moon-to-LLO link results in an outage for the scenario, the outage probability is expressed as:
Similarly, for the third scenario, where information is relayed through a GEO satellite, the outage probability is given below depending on the Moon-to-GEO and the GEO-to-Earth links.
The fourth scenario involves a multi-hop path: Moon-to-LLO, then LLO-to-GEO, and finally GEO-to-Earth. The outage probability is computed as:
4 DYNAMIC ORCHESTRATION: INTER-DOMAIN SPACE DIGITAL TWIN
As different CSN architectures provide varying advantages in terms of coverage and link reliability depending on dynamic operational conditions and mission requirements, a mechanism for intelligent and dynamic switching across lunar communication scenarios is critical. Therefore, we propose a digital twin that serves as an dynamic decision-making engine, modeling the entire communication system from lunar nodes to ground stations on Earth via LLO and GEO relays.
The primary goal of the digital twin is to optimize communication strategies during missions by balancing the crucial need for reliability with effective resource utilization, particularly in terms of transmit power, which improves overall system performance and enhances adaptability. Here, the inter-domain space digital twin is updated with the status, link conditions, and operational parameters of all relevant nodes each minute. Its comprehensive awareness allows to proactively assess and optimize for continuous communication, striving for stable reliability performance, and efficiency even when faced with potential risks or dynamic channel variations. This capability underpins its function of providing automated and informed decision making for network operations.
We present the core logic of the inter-domain digital twin via Algorithm 1 that hierarchically evaluates available communication scenarios by ensuring that the system always prioritizes achieving the best possible reliability. Initially, the algorithm assesses link conditions for all scenarios using the unified framework (Section 2) to compute their respective outage probabilities and required total transmit power. It then iterates through predefined reliability levels from highest to lowest. For each level, it identifies scenarios meeting the outage target and, crucially, selects the one that minimizes total transmit power. If no scenario meets even the low reliability requirement, a predefined fallback scenario is activated as baseline connectivity. In this way, it ensures reliable communication, optimizes power consumption, and automates complex operational decisions by dynamically selecting the most suitable scenario based on current conditions and mission priorities. We will further demonstrate the practical efficacy of this digital twin in Section 5.
5 NUMERICAL RESULTS
In this section, we present the numerical results obtained from simulations based on the framework and strategies described in Sections 2, 3 and 4. We first provide a detailed link analysis that is carried out on STK to ensure realistic orchestration of scenarios, as shown in Fig. 2. Then, we evaluate the reliability performance of each scenario and demonstrate the effectiveness of the digital twin algorithm to provide stable reliability.
cells = c,m,
row17 = c,row18 = c,row19 = c,
column5 = c,column6 = c,column7 = c,column8 = c,
cell11 = c=2c,cell13 = c,cell14 = c,
cell21 = c=2c,cell23 = c,cell24 = c,
cell31 = c=2c,cell34 = c=6,
cell41 = r=4c,cell42 = c,cell43 = c,cell44 = c,
cell54 = c, cell64 = c, cell74 = c,
cell81 = r=7c,cell82 = c,cell83 = c,cell84 = c,
cell94 = c, cell104 = c, cell114 = c,
cell124 = c, cell134 = c, cell144 = c,
cell151 = c=2c,
cell161 = c=2,cell164 = c=6,
cell171 = c=2c,
vlines, hline1-4,8,15-18 = -, hline5-7,9-15 = 2-9,
Parameters& Unit
Moon
to
Earth
Moon
to
LLO
LLO
to
Earth
Moon
to
GEO
GEO
to
Earth
LLO
to
GEO
GHz 26.25 27.25 26.25 27.25 26.25 27.25
MHz 50 50 50 50 50 50
Transmitter
W 2 0.5 2 2 0.01 2
0.85 0.85 0.90 0.85 0.93 0.90
m 1.5 1.5 1.5 1.5 4.6 1.5
dB 1.5 1.5 1 1.5 0.5 1
Receiver
0.95 0.90 0.95 0.93 0.95 0.93
m 34 1.5 34 4.6 34 4.6
dB 0.5 1 0.5 0.5 0.5 0.5
K 50 100 50 70 50 70
K 290 250 290 270 290 270
K 290 250 290 270 290 270
0.95 0.90 0.95 0.95 0.95 0.95
K 290 290 290
K 2.725 2.725 2.725 2.725 2.725 2.725
Rician- dB 10 20 15 13 30 40
5.1 Simulation Set-up and Configuration
The lunar communication architectures are simulated on STK through the four scenarios with a detailed link parameters in Table 1. The scenarios are evaluated minute-by-minute over a 24-hour period on October 1, 2024 with the following settings and configurations.
The Mons Malapert ()222The selenographic coordinate system is utilized to specify Mons Malapert’s location on the lunar surface., a promising candidate for future missions because of its potential water ice presence and near-consistent illumination [44], hosts the lunar communication terminal (LCT), as shown in Fig. 2.
Three DSN-NSN ground stations (GS) are located in Goldstone, California (GS-1); Canberra, Australia (GS-2); and Madrid, Spain (GS-3).
To maintain uninterrupted connectivity between the LLO network and the LCT, ten satellites are distributed equally across two orbital planes with inclinations of and .
In the beginning, LLO group 1 satellites are set with true anomalies of , whereas LLO group 2 satellites are assigned true anomalies of .
Six GEO relay satellites are deployed over the orbit, and the corresponding true anomaly values are set as , respectively.
Two-way communication is utilized by all relay satellites, which are equipped with antenna that can operate in both directions.
Since the Moon lacks an atmosphere like Earth’s and the lunar regolith reflects light differently than on Earth, we conduct our analyses by examining various illumination conditions of the lunar surface. These illumination fluctuations may also impact receiver antennas pointing toward an LLO satellite, exposing them to the lunar surface in the background. Therefore, the brightness temperature for antennas pointing to the Moon or the LLO satellite varied as , while it is fixed at for antennas pointing to GEO satellites.
5.2 In-Orbit Link Geometry and Propagation Dynamics
We analyze how geometrical factors of transmission dynamics—azimuth, elevation, and range (AER)—vary for communication links corresponding to four scenarios throughout a day, as shown in Fig. 3. These measurements are critical inputs for the unified link analysis framework, as they reveal significant variations in link conditions directly impacting communication visibility and signal propagation. Particularly, they highlight the challenges that necessitate inter-domain cooperation between space networks and their dynamic management.
We observe daily intervals of visibility and non-visibility for links in Fig. 3, as well as acquire insight into LoS conditions through changes in azimuth and elevation. The elevation angles of Moon-to-Earth direct links are usually around five degrees, denoting that the Moon is barely above the horizon from the perspective of three GS. Although there are still LoS, the fluctuating and low elevation means highly variable atmospheric path lengths for ground stations. When this is combined with ranges from roughly km to over km, the need for alternative relay solutions becomes clear. Moon-to-LLO links offer very short ranges, which are advantageous for minimizing path loss on the lunar access segment, but they have highly dynamic azimuth and elevation angles due to multiple LLO satellites. However, this benefit is countered by the rapid changes in azimuth and elevation, indicating a highly dynamic link and demanding competent antenna pointing and frequent handovers to maintain connectivity. LLO-to-Earth links exhibit higher azimuth and elevation compared to the Moon-to-Earth link, but at larger distances. These longer and highly variable distances, coupled with dynamic angles, mean significantly fluctuating path loss.
Moon-to-GEO links show smooth and periodic variations in AER that imply a more stable propagation environment than LLO links. The high distance of GEO relays allows for broad coverage of the cislunar space, and their consistent visibility provides opportunities for handovers to maintain connectivity; however, the considerable range of approximately km to km poses a challenge. In contrast, GEO-to-Earth links provide relatively stable azimuth and elevation over long periods with a short range around km. This means more predictable and lower free space path loss, which makes it more robust. LLO-to-GEO links (scenario 4) also show periodic but dynamic changes in AER. The complexity of such multi-hop links is evident, as they combine the characteristics of several individual links, inheriting their distinct dynamic behaviors. The pervasive dynamism among these potential links underscores the requirement for intelligent, adaptive, and dynamic management of these architectures to ensure continuous communication and use resources effectively while maximizing their advantages.
5.3 Architecture Reliability Assessment
This section examines the reliability performance of four lunar communication architectures with corresponding scenarios, building upon the dynamic link conditions established in Fig. 3. Fig. 4 illustrates the outage probability for each scenario over time under different brightness temperatures and an SNR threshold of 10 dB. Individual architecture analyses present their robustness for inter-domain space communication and implicitly lay the groundwork for the requirement of a dynamic decision-making mechanism among them.
The outage probability for direct Moon-to-Earth (scenario 1) communication exhibits significant variability and frequently reaches high values, as shown in Fig. 4a. This variability directly results from the visibility constraints observed in Fig. 3, where the elevations are very low, despite the presence of LoS via three ground stations. The high outage probabilities, particularly during high ranges, demonstrate that conventional direct communication is not sufficiently reliable for continuous operations on future lunar missions. The impact of increasing illumination is visible; while it influences the noise floor, the primary driver of outage in this architecture is the fundamental geometric and visibility limitation.
Scenario 2, utilizing LLO relays, shows improved reliability at first look in Fig. 4b due to the fewer colored points compared to direct communication. In fact, the number of high outages is larger, especially mostly around the times of handover between LLO satellites and with higher . This shows the dynamic nature of LLO links, as shown in Fig. 3, where rapid changes in visibility and frequent handovers might lead to instant changes for the links or outages. LLO relays offer lower path loss and better visibility in the cislunar domain (due to closer proximity to the lunar surface); however, the overall reliability depends on the availability and stability of link hops in the architecture. This means that maintaining a consistently low outage across both hops is challenging due to inherent orbital dynamics.
Scenario 3, relying on bidirectional GEO relay, generally demonstrates a more stable and often lower outage probability compared to scenario 1 and scenario 2. The outage probabilities mostly fall at and below, indicating significantly higher reliability. The smooth and periodic link dynamics of GEO satellites observed in Fig. 3 contribute to this improved stability. While higher values lead to an increase in the outage probability, the overall reliability is maintained at lower levels than other scenarios. Both results, in Fig. 3 and Fig. 4c, justify the advantages of the proposed NSN architecture for maintaining consistent and robust connectivity. over large distances.
The multi-hop scenario 4 generally exhibits the lowest outage probabilities among all scenarios across various values, except for the highest outages caused during the handover of LLO relay satellites as shown in Fig. 4d. This justifies that if this problem is overcome (such as by using more LLO satellites), the architecture will provide superior reliability by leveraging the benefits of both LLO (for short ranges and better visibility) and GEO for robust backhaul to Earth (stable and predictable link dynamics). On the other hand, boosting the multi-hop design with more satellites inherently means more redundancy, complexity, and resource usage. Therefore, the challenging trade-off between performance and cost constraints leads to scenario 3 being more amenable, while the higher performance demands make scenario 4 superior.
Across all scenarios, an increase in the brightness temperature typically correlates with an increase in outage probability. This is because contributes to the external antenna noise temperature, which in turn increases the operational equivalent noise temperature and thus the total noise power at the receiver. The sensitivity of outage probability to lunar surface brightness emphasizes the importance of accurately modeling the noise environment. Our comprehensive reliability analysis, showing the varied performance of each architecture, then establishes the requirement for dynamic management to maintain stable communication. This justifies the motivation for the development of the inter-domain space digital twin, as results show that no single architecture is always ideal and that a dynamically adaptive solution is necessary to overcome the inherent trade-offs between them.
5.4 Validation of Dynamic Orchestration
We observe the dynamic and complicated nature of lunar communication with significant variability in link geometries and challenges like limited line-of-sight, rapidly changing LLO relay links, and path losses via Fig. 3. Fig. 4 further reveals no single architecture consistently provides optimal reliability. This comprehensive analysis establishes that a dynamic orchestration is indispensable for continuous, robust, and resource-efficient lunar communication. Therefore, we introduce an inter-domain space digital twin that continuously models the entire communication system, dynamically optimizing strategies by balancing unwavering reliability with efficient transmit power utilization.
Given overall reliability performance with a relatively low requirement in Fig. 4, the third scenario—based on GEO relaying communication architecture—is set as a reserved fallback scenario. The practical efficacy of this digital twin is compellingly demonstrated in Fig. 5, which illustrates the reliability performance throughout a day under varying brightness temperatures. For a brightness temperature of 200 K, as shown in Fig. 5(a), the digital twin successfully maintains the communication link predominantly within the high reliability target (). The bar, showing chosen scenarios by minutes, indicates dynamic switching, primarily between scenario 3 and scenario 4, to achieve this consistent high reliability. This is crucial for mission success but also security, as it enables real-time command and control and vital data transmission, even under dynamic lunar conditions. The capability to achieve this while optimizing power consumption is pivotal, given that power is a scarce resource in the cislunar domain. In Fig. 5(a), we also observe occasional moderate reliability cases, but the system quickly recovers, and the fallback scenario (scenario 3) is rarely engaged. This highlights that our proposed inter-domain space digital twin can adapt to dynamic conditions (as in Fig. 3) and leverage architectural strengths for superior performance.
In Fig. 5(b), under an increased brightness temperature of 400 K, the digital twin continues to strive for high reliability. While largely operating within the high reliability level, more frequent and slightly longer periods show outage probability falling into moderate () or low reliability () levels. The fallback scenario is also utilized more often compared to Fig. 5(a) ( K). This demonstrates the digital twin’s robustness in adapting to more challenging noise environments, dynamically selecting scenarios to mitigate performance degradation. This dynamism is essential to overcome traditional communication limitations and ensures unwavering reliability, paramount for mission success and safety, even when maintaining the absolute highest reliability becomes more power intensive or geometrically constrained due to scarce power resources.
Overall, the inter-domain space digital twin provides a groundbreaking solution for advancing lunar communication. By continuously assessing link conditions, prioritizing unwavering reliability, and optimizing power consumption through dynamic scenario selection, it effectively overcomes the inherent limitations and complexities of individual architectures. This adaptive capability is fundamental for the success, safety, and scalability required for future lunar endeavors.
6 CONCLUSIONS
Establishing and maintaining reliable communication links is of paramount importance for lunar missions, as every phase of a mission, from telemetry and command to astronaut safety, depends on robust and uninterrupted connectivity. Failures or significant degradation in communication quality can jeopardize mission objectives and even endanger assets or human lives. This critical need for unwavering reliability motivates a move beyond conventional DTE communication, which is plagued by LoS limitations, economic non-scalability, and scheduling bottlenecks. This paper addresses these challenges by proposing a foundational shift towards inter-domain space network cooperation, integrating CSNs with NSNs.
To validate and compare these advanced architectures, we developed a unified link analysis framework that provides a high-fidelity assessment by incorporating unique environmental factors. Second, building on this framework, we assessed the reliability via outage probability to quantify the operational robustness of architectures under dynamic conditions. Through these, our work shows the advantages of NSN integrated architectures but also underscores the emerging requirement to build the resilient, dynamic, and high-performance communication backbone. Therefore, we proposed the inter-domain space digital twin, a dynamic decision-making engine that autonomously analyzes the networks in real-time to dynamically select the most reliable and power-efficient communication path. Final results show the proposed approach ensures optimal performance and resource utilization for the next generation of sustained lunar exploration.